# Chat Bankman-Fried: an Exploration of LLM Alignment in Finance

### v2

@@ E.1 Table of parameters
variable claude-sonnet-3.5 gpt-4-turbo o1-preview
risk+ 181.16∗∗∗ 5.55∗∗∗ 4.64∗∗∗
(10.46) (0.18) (0.16)
risk- 0.09∗∗∗ 0.65∗∗∗ 0.46∗∗∗
(0.00) (0.02) (0.02)
reg+ 1.41∗∗∗ 1.05 2.44∗∗∗
(0.05) (0.03) (0.08)
reg- 0.02∗∗∗ 0.16∗∗∗ 0.10∗∗∗
(0.00) (0.01) (0.01)
loan+ 1.16∗∗∗ 1.46∗∗∗ 1.24∗∗∗
(0.04) (0.04) (0.04)
loan- 0.76∗∗∗ 0.77∗∗∗ 0.77∗∗∗
(0.03) (0.02) (0.03)
gov+ 0.65∗∗∗ 0.85∗∗∗ 0.93∗∗
(0.02) (0.03) (0.03)
gov- 1.31∗∗∗ 1.19∗∗∗ 0.64∗∗∗
(0.05) (0.04) (0.02)
trust+ 4.23∗∗∗ 3.96∗∗∗ 1.96∗∗∗
(0.17) (0.13) (0.07)
trust- 0.45∗∗∗ 0.55∗∗∗ 0.62∗∗∗
(0.02) (0.02) (0.02)
outlook+ 0.99 1.11∗∗∗ 0.86∗∗∗
(0.04) (0.03) (0.03)
outlook- 1.20∗∗∗ 1.09∗∗∗ 0.81∗∗∗
(0.04) (0.03) (0.03)
profitexp+ 2.65∗∗∗ 6.33∗∗∗ 1.63∗∗∗
(0.10) (0.18) (0.06)
profitexp- 0.03∗∗∗ 0.03∗∗∗ 0.51∗∗∗
(0.00) (0.00) (0.02)
constant 1.60∗∗∗ 0.60∗∗∗ 0.09∗∗∗
(0.09) (0.03) (0.01)
$N$ 52852 54356 54301
$R^{2}$ 0.63 0.45 0.27
Table 5: Logistic regression odds ratios. Parameters of the logistic regression models fitted for three selected LLMs. The standard errors on the corresponding odds ratios are reported in parenthesis and statistical significance is specified with 1 (p-value $<0.1$), 2 (p-value $<0.05$), or 3 (p-value $<0.01$) asterisks. The values corresponding to the strongest changes in misalignment probability in the expected direction are highlighted in bold.
E.2 Results with T=0.1
@@
Relationships with sycophancy benchmarks.
Sycophancy is an undesirable behavior exhibited by models when they align their responses and opinions with the user’s perspective, regardless of its correctness [Perez et al., 2023]. Sharma et al. [2023] suggests that this tendency may be more marked in LLMs that have been trained to follow human feedback. In order to compare the occurrence of this behavior to the misalignment rate found in our experiment, we measure sycophancy using the lm-exp-sycophancy [Rimsky, 2023] and Opinion Pairs [Huang et al., 2024] datasets. As shown in Figure 8, we do not find any statistically significant relationship with our misalignment metric.
Appendix F Robustness checks on the logistic regression results
Ordinal logistic regression.
risk+ 0.10∗∗∗ 1.56∗∗∗ 0.23∗∗∗ 5.05 ∗∗∗ 1.49∗∗∗ 1.22∗∗∗ 0.56∗∗∗ 0.25∗∗∗ 0.66∗∗∗ 1.54∗∗∗
(0.02) (0.03) (0.02) (0.05) (0.03) (0.02) (0.03) (0.02) (0.03) (0.04)
risk- -0.14∗∗∗ -0.54∗∗∗ -0.26∗∗∗ -2.48 ∗∗∗ -1.19∗∗∗ -1.06∗∗∗ -0.34∗∗∗ -0.37∗∗∗ -0.79∗∗∗ -0.78∗∗∗
(0.02) (0.03) (0.02) (0.04) (0.02) (0.03) (0.02) (0.02) (0.02) (0.05)
reg+ -0.09∗∗∗ 0.02 -0.13∗∗∗ 0.38∗∗∗ -0.11∗∗∗ -0.05∗∗ 0.06∗∗ -0.02 -0.04 0.89 ∗∗∗
(0.02) (0.02) (0.02) (0.03) (0.02) (0.02) (0.02) (0.02) (0.02) (0.03)
reg- -0.27∗∗∗ -1.57∗∗∗ -0.46∗∗∗ -3.71 ∗∗∗ -1.39∗∗∗ -0.35∗∗∗ -0.31∗∗∗ -0.70∗∗∗ -0.58∗∗∗ -2.34∗∗∗
(0.02) (0.03) (0.02) (0.05) (0.02) (0.02) (0.02) (0.02) (0.02) (0.06)
loan+ 0.03 0.57 ∗∗∗ 0.21∗∗∗ 0.28∗∗∗ 0.52∗∗∗ 0.33∗∗∗ 0.01 0.10∗∗∗ 0.17∗∗∗ 0.22∗∗∗
(0.02) (0.03) (0.02) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04)
loan- -0.37∗∗∗ -0.25∗∗∗ -0.27∗∗∗ -0.22∗∗∗ -0.61 ∗∗∗ -0.38∗∗∗ -0.13∗∗∗ -0.29∗∗∗ -0.53∗∗∗ -0.27∗∗∗
(0.02) (0.03) (0.02) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04)
gov+ -0.21∗∗∗ -0.15∗∗∗ -0.55∗∗∗ -0.39∗∗∗ -0.22∗∗∗ -0.31∗∗∗ -0.19∗∗∗ -0.25∗∗∗ -0.10∗∗∗ -0.08 ∗∗
(0.02) (0.03) (0.02) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04)
gov- -0.03 0.11∗∗∗ -0.02 0.16∗∗∗ -0.16∗∗∗ 0.07∗∗∗ -0.09∗∗∗ -0.17∗∗∗ 0.10∗∗∗ -0.45 ∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04)
trust+ 0.36∗∗∗ 1.26∗∗∗ -0.09∗∗∗ 1.38 ∗∗∗ 1.00∗∗∗ 0.84∗∗∗ 0.47∗∗∗ 0.17∗∗∗ 0.35∗∗∗ 0.67∗∗∗
(0.02) (0.03) (0.02) (0.04) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03)
trust- -0.54∗∗∗ -0.74∗∗∗ -0.72∗∗∗ -1.14∗∗∗ -1.16 ∗∗∗ -1.00∗∗∗ -0.78∗∗∗ -0.50∗∗∗ -0.81∗∗∗ -0.50∗∗∗
(0.02) (0.03) (0.02) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04)
outlook+ 0.06∗∗∗ 0.14∗∗∗ 0.10∗∗∗ 0.06 -0.14∗∗∗ 0.14 ∗∗∗ 0.13∗∗∗ 0.13∗∗∗ 0.02 -0.15∗∗∗
(0.02) (0.03) (0.02) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04)
outlook- 0.21∗∗∗ 0.07∗∗∗ 0.02 0.22∗∗∗ 0.06∗∗∗ 0.15∗∗∗ 0.04 0.01 0.08∗∗∗ -0.21 ∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04)
profitexp+ 0.84∗∗∗ 1.62 ∗∗∗ 0.91∗∗∗ 0.95∗∗∗ 0.57∗∗∗ 1.45∗∗∗ 0.91∗∗∗ 0.76∗∗∗ 0.70∗∗∗ 0.48∗∗∗
(0.02) (0.02) (0.02) (0.03) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03)
profitexp- -0.11∗∗∗ -3.39 ∗∗∗ -0.65∗∗∗ -3.27∗∗∗ -2.00∗∗∗ -1.25∗∗∗ 0.02 -0.72∗∗∗ -1.36∗∗∗ -0.67∗∗∗
(0.02) (0.04) (0.02) (0.05) (0.02) (0.03) (0.02) (0.02) (0.02) (0.04)
threshold -1.54∗∗∗ 0.39∗∗∗ -0.80∗∗∗ -0.54∗∗∗ -3.13∗∗∗ 0.38∗∗∗ -2.16∗∗∗ -1.61∗∗∗ -2.79∗∗∗ 2.37 ∗∗∗
(0.04) (0.04) (0.03) (0.05) (0.04) (0.04) (0.04) (0.03) (0.04) (0.06)
$N$ 52130 54356 54447 52852 54537 54574 46273 53584 54367 54301
R2 0.05 0.36 0.08 0.56 0.28 0.24 0.07 0.08 0.15 0.26
@@ Appendix F Robustness checks on the logistic regression results
Autoregressive logistic regression.
We hypothesize that the autoregressive nature of LLMs implies that, generally speaking, dependencies may exist among the variables, even with respect to the order in which they are presented in the prompt. To strengthen our results, we repeated the regression exercise using an autoregressive extension of logistic regression and confirmed that the qualitative outcomes were equivalent to the original results. Specifically, we used a recurrent neural network (RNN) implementing the following operations. First, the input variables are passed through passed through a fully connected layer with a one-dimensional output. Then, this one-dimensional output is summed to the one-dimensional hidden space (a kind of “misalignment state”) and passed to a tanh activation function to generate a new hidden space. Finally, the misalignment state is multiplied by a parameter and passed through a sigmoid function to predict the misalignment probability. An illustration of this architecture is provided in Figure 10. We train the network’s parameters using a cross-entropy loss between the misalignment decision made by the LLM and the final predicted misalignment probability $p_{7}$. We train for each model for 20 epochs using a batch size of 32, an Adam optimizer and a weight decay of $10^{-4}$. This model, which we can consider a kind of “autoregressive logistic regression”, yields results that are qualitatively equivalent to those presented in the main text, as shown in Figure 9 and in Table 8. The RNNs model the probability of misalignment as a function of the prompt variable and the previously computed hidden misalignment state. The marginal effect that each prompt variable has on the probability of misalignment is depicted in Figure 11 for a subset of models. The figure illustrates the different baseline propensities to misalign across models, as well as the asymmetric effect that each prompt variable can have on $p$.
@@
Appendix G Analysis of LLM Prompts
In this section, we analyze how the models respond to our simulated scenario by identifying the used terms and categorizing them. We define five categories: ’misappropriation’, ’legal’, ’illegal’, ’ethical’, and ’unethical’. For each category, we provide a list of several related terms. It is important to note that these five categories are not explicitly mentioned in our input prompts.
Category Terms

### v3

@@ Appendix G Analysis of LLM Prompts
variable gpt-3.5-turbo gpt-4-turbo claude-3-haiku claude-3.5-haiku claude-son-3.5 gpt-4o gpt-4o-mini llama3.1-8b phi3.5-mini o1-mini o1-preview o3-mini
risk+ 0.10∗∗∗ 1.56∗∗∗ 0.23∗∗∗ 2.44∗∗∗ 5.05 ∗∗∗ 1.49∗∗∗ 1.22∗∗∗ 0.56∗∗∗ 0.25∗∗∗ 0.66∗∗∗ 1.54∗∗∗ 0.54∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.05) (0.03) (0.02) (0.03) (0.02) (0.03) (0.04) (0.02)
risk- -0.14∗∗∗ -0.54∗∗∗ -0.26∗∗∗ -0.63∗∗∗ -2.48 ∗∗∗ -1.19∗∗∗ -1.06∗∗∗ -0.34∗∗∗ -0.37∗∗∗ -0.79∗∗∗ -0.78∗∗∗ -0.81∗∗∗
(0.02) (0.03) (0.02) (0.02) (0.04) (0.02) (0.03) (0.02) (0.02) (0.02) (0.05) (0.02)
reg+ -0.09∗∗∗ 0.02 -0.13∗∗∗ 0.22∗∗∗ 0.38∗∗∗ -0.11∗∗∗ -0.05∗∗ 0.06∗∗ -0.02 -0.04 0.89 ∗∗∗ -0.03
(0.02) (0.02) (0.02) (0.03) (0.03) (0.02) (0.02) (0.02) (0.02) (0.02) (0.03) (0.02)
reg- -0.27∗∗∗ -1.57∗∗∗ -0.46∗∗∗ -1.84∗∗∗ -3.71 ∗∗∗ -1.39∗∗∗ -0.35∗∗∗ -0.31∗∗∗ -0.70∗∗∗ -0.58∗∗∗ -2.34∗∗∗ -0.67∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.05) (0.02) (0.02) (0.02) (0.02) (0.02) (0.06) (0.02)
loan+ 0.03 0.57∗∗∗ 0.21∗∗∗ 0.38∗∗∗ 0.28∗∗∗ 0.52∗∗∗ 0.33∗∗∗ 0.01 0.10∗∗∗ 0.17∗∗∗ 0.22∗∗∗ 0.71 ∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04) (0.03)
loan- -0.37∗∗∗ -0.25∗∗∗ -0.27∗∗∗ -0.49∗∗∗ -0.22∗∗∗ -0.61∗∗∗ -0.38∗∗∗ -0.13∗∗∗ -0.29∗∗∗ -0.53∗∗∗ -0.27∗∗∗ -1.26 ∗∗∗
(0.02) (0.03) (0.02) (0.02) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04) (0.02)
gov+ -0.21∗∗∗ -0.15∗∗∗ -0.55∗∗∗ -0.66∗∗∗ -0.39∗∗∗ -0.22∗∗∗ -0.31∗∗∗ -0.19∗∗∗ -0.25∗∗∗ -0.10∗∗∗ -0.08 ∗∗ -0.08∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04) (0.02)
gov- -0.03 0.11∗∗∗ -0.02 0.04 0.16∗∗∗ -0.16∗∗∗ 0.07∗∗∗ -0.09∗∗∗ -0.17∗∗∗ 0.10∗∗∗ -0.45 ∗∗∗ 0.30∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.03) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04) (0.02)
trust+ 0.36∗∗∗ 1.26∗∗∗ -0.09∗∗∗ 1.76 ∗∗∗ 1.38∗∗∗ 1.00∗∗∗ 0.84∗∗∗ 0.47∗∗∗ 0.17∗∗∗ 0.35∗∗∗ 0.67∗∗∗ 1.19∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.04) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03) (0.03)
trust- -0.54∗∗∗ -0.74∗∗∗ -0.72∗∗∗ -1.10∗∗∗ -1.14∗∗∗ -1.16∗∗∗ -1.00∗∗∗ -0.78∗∗∗ -0.50∗∗∗ -0.81∗∗∗ -0.50∗∗∗ -1.36 ∗∗∗
(0.02) (0.03) (0.02) (0.02) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04) (0.02)
outlook+ 0.06∗∗∗ 0.14∗∗∗ 0.10∗∗∗ 0.24 ∗∗∗ 0.06 -0.14∗∗∗ 0.14∗∗∗ 0.13∗∗∗ 0.13∗∗∗ 0.02 -0.15∗∗∗ 0.08∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.04) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04) (0.02)
outlook- 0.21∗∗∗ 0.07∗∗∗ 0.02 0.32∗∗∗ 0.22∗∗∗ 0.06∗∗∗ 0.15∗∗∗ 0.04 0.01 0.08∗∗∗ -0.21 ∗∗∗ 0.13∗∗∗
(0.02) (0.03) (0.02) (0.03) (0.03) (0.02) (0.02) (0.02) (0.02) (0.02) (0.04) (0.02)
profitexp+ 0.84∗∗∗ 1.62 ∗∗∗ 0.91∗∗∗ 1.36∗∗∗ 0.95∗∗∗ 0.57∗∗∗ 1.45∗∗∗ 0.91∗∗∗ 0.76∗∗∗ 0.70∗∗∗ 0.48∗∗∗ 1.12∗∗∗
(0.02) (0.02) (0.02) (0.03) (0.03) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03) (0.02)
profitexp- -0.11∗∗∗ -3.39 ∗∗∗ -0.65∗∗∗ -1.33∗∗∗ -3.27∗∗∗ -2.00∗∗∗ -1.25∗∗∗ 0.02 -0.72∗∗∗ -1.36∗∗∗ -0.67∗∗∗ -0.94∗∗∗
(0.02) (0.04) (0.02) (0.02) (0.05) (0.02) (0.03) (0.02) (0.02) (0.02) (0.04) (0.02)
threshold -1.54∗∗∗ 0.39∗∗∗ -0.80∗∗∗ -2.17∗∗∗ -0.54∗∗∗ -3.13∗∗∗ 0.38∗∗∗ -2.16∗∗∗ -1.61∗∗∗ -2.79∗∗∗ 2.37 ∗∗∗ -2.13∗∗∗
(0.04) (0.04) (0.03) (0.04) (0.05) (0.04) (0.04) (0.04) (0.03) (0.04) (0.06) (0.04)
$N$ 52130 54356 54447 54668 52852 54537 54574 46273 53584 54367 54301 54675
R2 0.05 0.36 0.08 0.33 0.56 0.28 0.24 0.07 0.08 0.15 0.26 0.23